OpenStudy (anonymous):

I have the equation I just need help figuring it out! Please! 3.Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+r /n )^1 /c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+r/n. My equation- 10,000=39,145[(1+0.03/12)^1/4]^(4)12t

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

OpenStudy (anonymous):

@Hero @ParthKohli

OpenStudy (anonymous):

@texaschic101 @SolomonZelman

hero (hero):

Lastly? What happened Firstly?

OpenStudy (anonymous):

There was two other questions but I answered them already I'm just having a hard time with this one and need to know how to figure out that equation. @Hero

hero (hero):

The questions all go together. It is not possible to answer the third question without knowing what the first two are about. It's like giving someone the end of a novel and expecting them to know what happened the beginning.

OpenStudy (anonymous):

1.Using the function A(t)=P(1+r/ n )nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. 2. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. 3.Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+r /n )1 /c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+r/ n. @hero

OpenStudy (anonymous):

My answer for 1. is A(t) = 39,145(1+ 0.03/12)^(12t) my answer for 2. is -45.5=t

hero (hero):

Where did you get the information for the first question?

OpenStudy (anonymous):

It tells me in my lesson what each letter represents. I really just need help solving the equation that's all please. @hero

hero (hero):

You mean you need help solving this: 10,000=39,145[(1+0.03/12)^1/4]^(4)12t

OpenStudy (anonymous):

Yes.

hero (hero):

\[10,000=39,145\left[\left(1+\frac{0.03}{12}\right)^{\frac{1}{4}}\right]^{(4)12t}\]

hero (hero):

Basically, you are solving for t. I can walk you through the steps. For each step I suggest for you to do, post the result you get afterwards, so I can keep track.

OpenStudy (anonymous):

Okay thank you so much!

hero (hero):

Well, technically you could just take logs of both sides. If you had a calculator, you could just log both sides, then isolate t.

OpenStudy (anonymous):

Yeah but wouldn't I have simplify all the exponents?

hero (hero):

Not necessarily. If you have a calculator, all you have to do is just input the information into your calculator correctly. Let's try to do this step by step. First log both sides. Let me know what you get for the result.

hero (hero):

Remember this rule \[\log(a^b) = b \log(a)\]

OpenStudy (anonymous):

log of 10,000 would be 4 correct?

hero (hero):

Yeah, but you didn't have to perform the log just yet. But that at least simplifies the left side.

hero (hero):

Once you have logged both sides, write out what you think the equation looks like afterwards.

OpenStudy (anonymous):

10,000=39,145 (1.0025^1/4)^48t

OpenStudy (anonymous):

right?

hero (hero):

Log both sides. You had 4 for the left side after logging right? Now log the other side. Let me know what you get afterwards.

OpenStudy (anonymous):

but I don't get how I can log it if it has the 1/4?

hero (hero):

Don't worry about the 1/4 just yet. Just apply the rule \(\log(a^b) = b \log(a)\) for the 48t

OpenStudy (anonymous):

Oh! so it'd be 48t log(1.0025)^1/4 right?

hero (hero):

On the right side yes

hero (hero):

Well wait, the 1/4 should be inside the parentheses

hero (hero):

You should have 4 = 48t log(1.0025^{1/4})

hero (hero):

Now all you have to do is just isolate t

OpenStudy (anonymous):

so then itd be 4/48 log (1.0025^(1/4))=t?

hero (hero):

yes, but the value you get for t won't mean anything

OpenStudy (anonymous):

so now what?

hero (hero):

If I were you I'd double check your values to make sure you have them correct.

hero (hero):

Does it say what t represents in this case?

OpenStudy (anonymous):

I have. :( im so tired of this ive been working on this stupid question for two hours already im just going to call my teacher.

hero (hero):

Then it is saying that t will be negative.

hero (hero):

Even stranger

OpenStudy (anonymous):

Wait yes it is supposed to be a negative

hero (hero):

Actually, it says to solve for the adjusted interest rate.

hero (hero):

For part 3

hero (hero):

So looks like you set something up wrong to begin with. In step 2, you should have already solved for t it seems.

OpenStudy (anonymous):

I did t=-45.5

hero (hero):

It seems to me that in step 3, you should be solving for \(r\) , not \(t\)

hero (hero):

Where did you get the 0.03 from?

OpenStudy (anonymous):

the percent interest rate.

hero (hero):

It says to solve for the adjusted interest rate.

hero (hero):

You might need to call your teacher for this.

hero (hero):

The problem is you have values that I don't even know where you got them.

OpenStudy (anonymous):

I just want to cry im so frustrated and I want to give up.

OpenStudy (anonymous):

it gave me them First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3%

hero (hero):

@Mollyyy, just a suggestion... When posting a question, post all of the relevant information. You shouldn't assume that you did anything right. The only thing you should do is post all of the information related to the problem and then post what you have done so far. The first thing you told me was that all you had to do was just solve the equation for t. Once we solved it for t, I immediately knew something was wrong. But the whole time I was thinking to myself, why does this user assume she has everything correct so far and why hasn't she posted all of the information necessary to solve the problem?

hero (hero):

I understand now why you want to cry.

OpenStudy (anonymous):

Well there's all the information. :(

hero (hero):

Yeah, but only the information for step 3.

hero (hero):

You still haven't answered my question about where you got the information for steps 1 and 2.

hero (hero):

Trust me, at this point, you are better off getting help from your teacher.

OpenStudy (anonymous):

:( okay thanks. and sorry for wasting all this time of yours for nothing.